OpenStudy (anonymous):

Please help. Will Medal. Given the following functions f(x) and g(x), solve for (f ⋅ g)(2) and select the correct answer below: f(x) = 3x2 + 2 g(x) = x − 8 −42 −84 42 84

OpenStudy (anonymous):

@zepdrix @Zarkon @acxbox22 @Awolflover1 @Aureyliant @dan815 @Data_LG2

OpenStudy (anonymous):

first just plug in 2 as x in both of the formulas f(x) = 3(2)^2 + 2 g(x) = (2) − 8

OpenStudy (anonymous):

solve that and multiply the numbers you get times each other